### v1

@@ Abstract
Explainable artificial intelligence (XAI) has become increasingly important in decision-critical domains such as healthcare, finance, and law. Counterfactual (CF) explanations, a key approach in XAI, provide users with actionable insights by suggesting minimal modifications to input features that lead to different model outcomes. Despite significant advancements, existing CF generation methods often struggle to balance proximity, diversity, and robustness, limiting their real-world applicability. A widely adopted framework, Diverse Counterfactual Explanations (DiCE), emphasizes diversity but lacks robustness, making CF explanations sensitive to perturbations and domain constraints. To address these challenges, we introduce DiCE-Extended, an enhanced CF explanation framework that integrates multi-objective optimization techniques to improve robustness while maintaining interpretability. Our approach introduces a novel robustness metric based on the Dice-Sørensen coefficient, ensuring stability under small input variations. Additionally, we refine CF generation using weighted loss components ($\lambda_{p}$, $\lambda_{d}$, $\lambda_{r}$) to balance proximity, diversity, and robustness. We empirically validate DiCE-Extended on benchmark datasets (COMPAS, Lending Club, German Credit, Adult Income) across multiple ML backends (Scikit-learn, PyTorch, TensorFlow). Results demonstrate improved CF validity, stability, and alignment with decision boundaries compared to standard DiCE-generated explanations. Our findings highlight the potential of DiCE-Extended in generating more reliable and interpretable CFs for high-stakes applications. Future work will explore adaptive optimization techniques and domain-specific constraints to further enhance CF generation in real-world scenarios.
@@ 1 Introduction
To address these challenges, we propose DiCE-Extended, an advanced framework designed to enhance the robustness of CF generation. Our contributions are as follows:
  1. 1.
  2. 2.
  3. 3.
    Empirical Machine Learning (ML) Validation: We evaluate DiCE-Extended on benchmark datasets (COMPAS, Lending Club, German Credit, Adult Income), demonstrating controllable performance over the original DiCE framework in fidelity, stability, and practical applicability [33], [40], [23].
The remainder of this paper is structured as follows: Section 2 discusses the theoretical foundations of CF generation and its associated challenges. Section 3 presents the methodology behind DiCE-Extended, including its optimization framework and robustness integration. Section 4 details experimental evaluations and performance comparisons. Finally, Section 5 concludes the study and outlines future research directions.
2 Background
The increasing adoption of ML in high-stakes domains such as healthcare, finance, and criminal justice necessitates transparent and interpretable models. However, complex models often function as “black boxes”, limiting trust in their predictions [18], [17]. XAI addresses this challenge through interpretability techniques like SHAP and LIME, which provide feature-based explanations but sometimes compromise fidelity [2], [28], [39]. Feature importance methods, including permutation importance, improve transparency by mitigating biases in variable relevance [1]. Additionally, tree-based models benefit from game-theoretic approaches for enhanced local and global interpretability [32]. Despite these advances, ensuring actionable and domain-relevant explanations remains a challenge, requiring integration of causal reasoning and balancing trade-offs between simplicity and accuracy [7], [38].
2.1 Approaches to Counterfactual Explanations
CF explanations provide “what-if” scenarios by suggesting minimal modifications to input features that alter a model’s prediction. Traditional CF explanations methods focus on perturbation vectors but often overlook feature dependencies and practical feasibility. Decision-tree-based techniques like CF Explanation Trees [25] improve interpretability, while Ordered CF explanations [26] prioritize feature modifications based on asymmetric dependencies. Distribution-aware methods such as DiCE enhance plausibility by incorporating feature correlations [24]. Cutting-edge approaches such as ReLAX leverage deep reinforcement learning for efficient and scalable counterfactual generation [5], [6], while interactive platforms like ViCE enhance user engagement through intuitive visualization tools [19]. However, despite these advancements, CF explanations remain susceptible to adversarial manipulation, highlighting the need for robust mechanisms that ensure reliability, fairness, and trust [4], [30].
2.2 Ensuring Robustness in Counterfactual Generation
Ensuring robustness in CF explanations is crucial for their reliability and practical application. Bansal et al. [3] demonstrated that robust training enhances the consistency of explanations, reducing sensitivity to model variations. Similarly, Datta et al. [9] highlighted that mitigating adversarial vulnerabilities strengthens the interpretability of CF explanations, making them more resistant to manipulation. To further enhance robustness, Huai et al. [huai2022automating] introduced an automated concept-based framework that provides certified robustness guarantees, minimizing the need for human intervention. Virgolin and Fracaros [36] proposed incorporating robustness constraints into CF generation, ensuring reliability even under input perturbations. Additionally, Bayesian frameworks for uncertainty quantification, as explored by Slack et al. [34], improve the stability of CF explanations by accounting for uncertainty in predictions. Finally, Dombrowski et al. [11] emphasized the importance of generating plausible CF explanations, which enhances fairness and ensures that suggested changes remain feasible within real-world constraints.
2.3 Challenges and Open Problems
Despite progress, CFEs face several persistent challenges, including robustness, actionability, scalability, and fairness, as summarized in Table 1.
@@ 3.1.1 Objective:
Given an input instance $\mathbf{x}\in\mathbb{R}^{d}$ and a trained ML model $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$, the task is to generate a set of $k$ counterfactual examples ${\mathbf{c}_{1},\mathbf{c}_{2},\dots,\mathbf{c}_{k}}$ that influence the model’s decision toward a specified target behavior while maintaining interpretability and actionability.
3.1.2 Input Components
@@ 3.2 Hypotheses
The generation of CF explanations follows these hypotheses:
@@ 3.3 Mathematical Formalizations
$\begin{split}\mathcal{L}(\mathbf{c}_{1},\dots,\mathbf{c}_{k})=\underset{c_{1},...,c_{k}}{\text{arg min}}[{\mathcal{L}_{\text{y\_loss}}}+\lambda_{p}\mathcal{L}_{\text{Proximity}}-\\ \lambda_{d}\mathcal{L}_{\text{Diversity}}-\lambda_{r}\mathcal{L}_{\text{Robustness}}],\end{split}$ (1)
where $\lambda_{p}$, $\lambda_{d}$, and $\lambda_{r}$ are hyperparameters balancing proximity, diversity, and robustness, respectively.
@@ 3.3.1 Proximity Metric:
The proximity metric ensures minimal deviation from $\mathbf{x}$, computed as:
$\text{Proximity}=\frac{1}{k}\sum_{i=1}^{k}|\mathbf{c}_{i}-\mathbf{x}|_{p}.$ (2)
$\|\cdot\|_{p}$ represents a norm-based distance metric, such as:
  • $\ell_{1}$ norm: $\|\mathbf{c}_{i}-\mathbf{x}\|_{1}=\sum_{j=1}^{d}|c_{i,j}-x_{j}|$.
  • $\ell_{2}$ norm: $\|\mathbf{c}_{i}-\mathbf{x}\|_{2}=\sqrt{\sum_{j=1}^{d}(c_{i,j}-x_{j})^{2}}$.
Proximity can be incorporated into the objective function as a penalty term to minimize unnecessary deviations from the original input:
$\mathcal{L}_{\text{Proximity}}=\lambda_{p}\cdot\|\mathbf{c}_{i}-\mathbf{x}\|_{p},$ (3)
where $\lambda_{p}$ is a hyperparameter that controls the trade-off between proximity and other objectives, set to 0.5 as specified in the DiCE framework [33].
@@ 3.3.2 Diversity Metric:
Diversity among CF explanations is measured via a similarity kernel matrix $\mathbf{K}$:
$\text{Diversity}=\det(\mathbf{K}).$ (4)
where $\mathbf{K}$ is a similarity kernel matrix defined as:
$\mathbf{K}_{i,j}=\frac{1}{1+\text{dist}(\mathbf{c}_{i},\mathbf{c}_{j})},$ (5)
and $\text{dist}(\cdot,\cdot)$ is a distance metric, such as $\ell_{1}$ or $\ell_{2}$ norm, to measure the difference between CF explanations $\mathbf{c}_{i}$ and $\mathbf{c}_{j}$.
Diversity can be explicitly incorporated into the loss function to encourage variability among CF explanations:
$\mathcal{L}_{\text{Diversity}}=-\lambda_{d}\cdot\det(\mathbf{K}),$ (6)
where $\lambda_{d}$ is a hyperparameter that regulates the trade-off between diversity and other objectives. A higher determinant value indicates a more diverse set of CF explanations. Following the DiCE framework [33], $\lambda_{d}=1.0$ is set to 1.0.
@@ 3.3.3 Robustness Metric:
Robustness is computed as:
$\text{Robustness}=\frac{2\cdot|\mathbf{c}\cap\mathbf{c}^{\prime}|}{|\mathbf{c}|+|\mathbf{c}^{\prime}|},$ (7)
where $\mathbf{c}$ is CF explanations generated for $\mathbf{x}$ and $\mathbf{c}^{\prime}$ is perturbed CF explanations as $\mathbf{c}^{\prime}=\mathbf{c}+\boldsymbol{\delta}$, and $\boldsymbol{\delta}$ represents a slight noise, respectively.
During the robustness loss computation, counterfactuals and perturbed counterfactuals are binarized to ensure discreteness, then the resulting binary vectors are sent to Dice-Sørensen coefficient function. Robustness can be enhanced by introducing a regularization term into the objective function:
@@
where $\lambda_{r}$ is a hyperparameter, and $robustness\_loss$ measures the distance between CF explanations with Dice-Sørensen coefficient.
We performed a grid search for the robustness weight $\lambda_{r}$ within the range $[0,1]$. As shown in Figure 1, the robustness loss reaches its peak at $\lambda_{r}=0.4$, while the total loss continues to decrease beyond this value. Additionally, the variation in computed robustness across the grid is minimal, approximately $3\times 10^{-4}$. To maintain a balanced trade-off between proximity, diversity, and robustness, we selected $\lambda_{r}=0.4$ as a moderate value. This ensures that robustness remains a contributing factor to the overall loss without dominating the optimization process.
Figure 1: Impact of robustness weight ($\lambda_{r}$) on robustness loss and total loss.
3.3.4 Constraints:
CF explanations must adhere to the following constraints:
  • Feature Constraints: Each feature value must lie within its valid range, specified by $\mathcal{C}$.
  • Actionability Constraints: Certain features may be immutable (e.g., gender or age) and must not be altered in the CF.
@@ 3.4 Dataset Information
To ensure the robustness of our CF explanation framework, we evaluate it on four diverse datasets [33].
  • Adult-Income: Extracted from the UCI Machine Learning Repository, this dataset includes demographic and employment-related features. We use 8 key attributes, such as hours per week, education level, and occupation, to classify individuals earning above or below \$50,000 per year.
  • LendingClub: This dataset contains five years of loan records (2007-2011) from LendingClub. We select 8 essential features, including employment years, annual income, and credit history, to predict whether a borrower will repay their loan.
  • German-Credit: This dataset includes 20 financial and demographic attributes related to loan applicants. The task is to classify applicants as good or bad credit risks.
  • COMPAS: A dataset from ProPublica focused on recidivism prediction in the U.S. criminal justice system. After preprocessing, 5 key features are retained (e.g., age, prior offenses, and charge severity) to predict whether a defendant will reoffend within two years.
3.5 Experimental Setup
To systematically evaluate the performance of the proposed CF explanation framework, we conduct experiments on all four datasets using three distinct ML frameworks: Scikit-learn, PyTorch (PYT), and TensorFlow (TF2). The evaluation follows a structured approach, ensuring fair comparisons across different models and loss function components.
@@ 3.5.1 Model Architectures
The experimental setup includes two distinct model architectures, each designed to assess CF explanations from different learning paradigms:
@@
where $W_{1}\in\mathbb{R}^{n\times d}$ and $W_{2}\in\mathbb{R}^{1\times n}$ are the weight matrices where $n$ is number of features, $b_{1}\in\mathbb{R}^{n}$ and $b_{2}\in\mathbb{R}$ are bias terms, and $\sigma$ represents the Sigmoid activation function. This neural network is trained via backpropagation using a binary cross-entropy loss function. The chosen architecture ensures model flexibility while maintaining computational efficiency during CF generation.
@@ 3.5.2 Training Configuration
To ensure consistency in model evaluation, we adhere to a standardized training protocol across all frameworks. The training process incorporates the following hyperparameter configurations:
  • Learning Rate: Set to 0.001, ensuring stable gradient updates while mitigating convergence issues.
@@
To prevent overfitting, early stopping mechanisms are integrated, monitoring validation loss to halt training if no improvement is observed. Additionally, model evaluation metrics, including accuracy and loss progression, are continuously logged to assess the stability of CF explanations.
@@ 4.1 Loss Progression in CF Explainers
We analyzed the evolution of different loss components—including robustness loss—during the CF generation process across four datasets. Figure 2 illustrates how the optimization process minimizes these losses over multiple iterations, ensuring stable and interpretable CF explanations.
Focusing on robustness loss metric, we observe an initial fluctuation in the early iterations, followed by a steady decline as the optimization progresses. By approximately the 50th iteration, robustness loss stabilizes, indicating that the generated CFs maintain consistency despite small perturbations in input features. This trend suggests that the optimization framework successfully enhances the resilience of CF explanations, making them more reliable for real-world applications. The convergence of robustness loss aligns with the overall loss minimization strategy, balancing proximity, diversity, and stability without compromising interpretability. Moreover, the stabilization of robustness loss demonstrates that after a certain number of iterations, further optimization does not significantly improve CF reliability. This finding highlights the efficiency of the framework in generating CFs within a limited number of steps while ensuring robustness in decision-making scenarios.
4.2 Performance Comparison: Original vs. DiCE vs. DiCE-Extended Scenarios
To evaluate the effectiveness of DiCE-Extended, we compare ML model performance across three conditions: (1) models trained and tested on the original datasets, (2) models trained on original datasets but tested on CF datasets generated by DiCE, and (3) models trained on original datasets but tested on CF datasets generated by DiCE-Extended. The comparison focuses on key classification metrics, including accuracy, F1-score, recall, precision, AUC, as well as generation time and validity scores of CF samples.
4.2.1 Key Observations and Findings
1. Original Model Performance (Baseline Evaluation):
Models trained and tested on the original datasets generally exhibit high accuracy, particularly in datasets like Adult Income and Lending Club, where scores exceed 80% accuracy in most cases (Table 2). However, performance varies across different backends (Scikit-learn, PYT, TF2) due to their respective architectures and optimization mechanisms.
2. DiCE CFs (Baseline CF Evaluation):
When models trained on the original datasets were tested on CF examples generated by DiCE, there was a notable performance degradation. While Scikit-learn models retained perfect accuracy across datasets, DL models in PYT and TF2 showed significant drops in accuracy and F1-score, particularly in high-dimensional datasets like German Credit (Table 3). This suggests that standard DiCE-generated CFs may not always align with the decision boundaries of DL models, impacting model validity.
3. DiCE-Extended CFs (Optimized CF Evaluation):
The introduction of DiCE-Extended resulted in more stable CF explanations, leading to improved validity and classification performance compared to the original DiCE framework (Table 4). Notably:
  • PYT models showed improved robustness, with higher F1-scores and AUC values compared to their performance with standard DiCE CFs.
  • TF2 models demonstrated increased validity, particularly in high-dimensional datasets like COMPAS-Recidivism, where DiCE-Extended CFs aligned better with the decision function.
  • Generation time increased for TF2-based models, reflecting the additional optimization complexity introduced by robustness constraints in DiCE-Extended.

### v2

@@ Abstract
Explainable artificial intelligence (XAI) has become increasingly important in decision-critical domains such as healthcare, finance, and law. Counterfactual (CF) explanations, a key approach in XAI, provide users with actionable insights by suggesting minimal modifications to input features that lead to different model outcomes. Despite significant advancements, existing CF generation methods often struggle to balance proximity, diversity, and robustness, limiting their real-world applicability. A widely adopted framework, Diverse Counterfactual Explanations (DiCE), emphasizes diversity but lacks robustness, making CF explanations sensitive to perturbations and domain constraints. To address these challenges, we introduce DiCE-Extended, an enhanced CF explanation framework that integrates multi-objective optimization techniques to improve robustness while maintaining interpretability. Our approach introduces a novel robustness metric based on the Dice-Sørensen coefficient, enabling stability under small input variations. Additionally, we refine CF generation using weighted loss components (${\lambda_{p}}$, ${\lambda_{d}}$, ${\lambda_{r}}$) to balance proximity, diversity, and robustness. We empirically validate DiCE-Extended on benchmark datasets (COMPAS, Lending Club, German Credit, Adult Income) across multiple ML backends (Scikit-learn, PyTorch, TensorFlow). Results demonstrate improved CF validity, stability, and alignment with decision boundaries compared to standard DiCE-generated explanations. Our findings highlight the potential of DiCE-Extended in generating more reliable and interpretable CFs for high-stakes applications. Future work could explore adaptive optimization techniques and domain-specific constraints to further enhance CF generation in real-world scenarios.
@@ 1 Introduction
However, there is still place for improvement in CF explanation methods as the summary of challenges in CF explanations identified in the literature review is provided in Table 1.
@@
To tackle the challenges outlined in Table 1, we propose DiCE-Extended, an advanced framework designed to enhance the robustness of CF generation. Our contributions are as follows:
  1. 1.
  2. 2.
  3. 3.
    Empirical Validation on Diverse Benchmarks: We benchmark DiCE-Extended on four widely used datasets (Adult-Income, Lending-Club, German-Credit, and COMPAS) under three implementation back-ends (Scikit-learn, PyTorch, TensorFlow). Across this setup, the new framework consistently equals or surpasses the original DiCE: proximity decreases substantially, yet validity stays near 100%; both sparsity and diversity rise; and robustness improves markedly for the DL models while remaining at least slightly higher for tree-based pipelines. Finally, the 1-NN fidelity is preserved and often improved, indicating that the observed gains do not compromise the integrity of the local decision boundary.
The remainder of this paper is structured as follows: Section 2 presents the methodology behind DiCE-Extended, including its optimization framework and robustness integration. Section 3 details experimental evaluations and performance comparisons. Finally, Section 4 concludes the study and outlines future research directions.
@@ 2.1.1 Objective:
Given an input instance $\mathbf{x}\in\mathbb{R}^{d}$ and a trained ML model $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$, the task is to generate a set of $k$ CF examples ${\mathbf{c}_{1},\mathbf{c}_{2},\dots,\mathbf{c}_{k}}$ that influence the model’s decision toward a specified target behavior while maintaining interpretability and actionability.
2.1.2 Input Components:
@@ 2.2 Hypotheses
The generation of CF explanations follows the hypotheses:
@@ 2.3 Mathematical Formalizations
$\mathcal{L}(\mathbf{c})=\operatorname*{arg\,min}_{c_{1},\ldots,c_{k}}\bigl{(}\mathcal{L}_{y\_{\text{loss}}}+\lambda_{p}\mathcal{L}_{\text{Proximity}}-\lambda_{d}\mathcal{L}_{\text{Diversity}}-\lambda_{r}\mathcal{L}_{\text{Robustness}}\bigr{)},$
where $\lambda_{p}$, $\lambda_{d}$, and $\lambda_{r}$ are hyperparameters that balance proximity, diversity, and robustness, respectively.
@@ 2.3.1 Proximity Metric:
The proximity metric enables minimal deviation from $\mathbf{x}$, computed by $\frac{1}{k}\sum_{i=1}^{k}|\mathbf{c}_{i}-\mathbf{x}|_{p}$ which is a norm-based distance metric. Proximity can be incorporated into the objective function as a penalty term to minimize unnecessary deviations from the original input with $\mathcal{L}_{\text{Proximity}}=\lambda_{p}\cdot\|\mathbf{c}_{i}-\mathbf{x}\|_{p},$ where $\lambda_{p}$ is a hyperparameter that controls the trade-off between proximity and other objectives. Here, $\lambda_{p}$ is set to 0.5, matching the value used in the original DiCE framework to enable a fair comparison with the benchmark method DiCE [19].
2.3.2 Sparsity Metric:
The sparsity metric quantifies the number of features that are needed to alter the output of a black-box model. In this study, the number of different features between the input instance and the generated CF is measured using this metric [19]. The sparsity is defined as the average of the number of different features deducted from 1 which can be written mathematically by $1-\frac{1}{k\times d}\,\sum_{i=1}^{k}\sum_{j=1}^{d}\mathbf{1}\!\left[c_{i}^{\,j}\neq x^{\,j}\right]$.
@@ 2.3.3 Diversity Metric:
Diversity among CF explanations is defined with $\det(\mathbf{K})$ where $\mathbf{K}$ refers to a similarity kernel matrix. The kernel matrix is defined by $\mathbf{K}_{i,j}=1/[{1+\text{dist}\,(\mathbf{c}_{i},\mathbf{c}_{j})]}$, where $\text{dist}(\cdot,\cdot)$ is a distance metric such as $\ell_{1}$ or $\ell_{2}$ norm to measure the difference between CF explanations $\mathbf{c}_{i}$ and $\mathbf{c}_{j}$.
Diversity can be explicitly incorporated into the loss function to encourage variability among CF explanations which can be formulated by $\mathcal{L}_{\text{Diversity}}=-\lambda_{d}\cdot\det(\mathbf{K})$, where $\lambda_{d}$ is a hyperparameter that regulates the trade-off between diversity and other objectives. The higher the determinant value is, there are more diverse the set of CF explanations. Following the DiCE framework [19], $\lambda_{d}=1.0$ is set to 1.0.
@@ 2.3.4 Robustness Metric:
To compute the robustness metric, we adopt the Dice–Sørensen coefficient, defined as $\frac{2\cdot|\mathbf{c}\cap\mathbf{c}^{\prime}|}{|\mathbf{c}|+|\mathbf{c}^{\prime}|}$. In this context, $\mathbf{c}$ denotes the CF explanations generated for the input $\mathbf{x}$, while $\mathbf{c}^{\prime}$ represents the perturbed CF explanations, given by $\mathbf{c}^{\prime}=\mathbf{c}+\boldsymbol{\delta}$, where $\boldsymbol{\delta}$ denotes the perturbation applied to the generated CFs.
During the computation of the robustness loss, both the CFs and their perturbed counterparts are binarized to ensure discrete representations. The Dice–Sørensen coefficient is then used to measure the distance between the resulting binary vectors. Robustness is incorporated into the overall objective function through a regularization term defined as:
@@
where $\lambda_{r}$ is a tunable hyperparameter and robustness_loss corresponds to the Dice–Sørensen distance between the original and perturbed CF explanations.
To determine an appropriate value for $\lambda_{r}$, we conducted a grid search within the range $[0,1]$ across all datasets. For example, on the Adult Income dataset using the PyTorch backend, the robustness loss was found to peak at $\lambda_{r}=0.4$, although the total loss continued to decrease beyond this point. Moreover, the variation in the computed robustness across the grid was relatively small, approximately $3\times 10^{-4}$. Based on these findings, we selected $\lambda_{r}=0.4$ as a moderate and balanced value, ensuring that robustness contributes meaningfully to the overall optimization without dominating the objectives of proximity and diversity. The same grid search procedure was applied to all datasets.
@@ 2.4 Dataset Information
To assess the robustness of our CF explanation framework, we conduct evaluations across four diverse benchmark datasets.
  • Adult-Income [2]: Extracted from the UCI Machine Learning Repository, this dataset includes demographic and employment-related features. We use 8 key attributes, such as hours per week, education level, and occupation, to classify individuals earning above or below \$50,000 per year.
  • LendingClub [1]: This dataset contains five years of loan records (2007-2011) from LendingClub. We select 8 essential features, including employment years, annual income, and credit history, to predict whether a borrower will repay their loan.
  • German-Credit [15]: This dataset includes 20 financial and demographic attributes related to loan applicants. The task is to classify applicants as good or bad credit risks.
  • COMPAS [11]: A dataset from ProPublica focused on recidivism prediction in the U.S. criminal justice system. After preprocessing, 5 key features are retained (e.g., age, prior offenses, and charge severity) to predict whether a defendant will reoffend within two years.
2.5 Experimental Setup for Robustness Assessment
To benchmark DiCE-Extended against the original DiCE explainer, we implemented a unified evaluation pipeline across four tabular datasets (Adult-Income, Lending-Club, German-Credit, and COMPAS-Recidivism) and three model back-ends: Scikit-learn, PyTorch, and TensorFlow.
@@ 2.5.1 Model Architectures
The experimental setup incorporates two distinct model architectures, each tailored to evaluate CF explanations from different learning paradigms:
@@
where $W_{1}\in\mathbb{R}^{n\times d}$ and $W_{2}\in\mathbb{R}^{1\times n}$ are the weight matrices. Here $n$ is the number of features, $b_{1}\in\mathbb{R}^{n}$ and $b_{2}\in\mathbb{R}$ are bias terms, and $\sigma$ represents the Sigmoid activation function. This neural network is trained via backpropagation using a binary cross-entropy loss function. The chosen architecture enables model flexibility while maintaining computational efficiency during CF generation.
@@ 2.5.2 Training Configuration
For consistent model evaluation, we follow a standardized training protocol across all back-end frameworks. The training process employs the following hyperparameter configurations:
  • Learning Rate: Set to 0.001 to promote stable gradient updates and mitigate convergence issues.
@@
To prevent overfitting, early stopping mechanisms are employed by monitoring the validation loss and terminating training when no further improvement is observed. Additionally, evaluation metrics such as accuracy and loss progression are continuously logged to track model performance and assess the stability of CF explanations.
@@ 3.1 Loss Progression in CF Explainers
We analyzed the evolution of various loss components, including robustness loss, during the CF generation process across four datasets. Figure 1 illustrates how the optimization procedure reduces these losses over successive iterations, leading to stable and interpretable CF explanations. Focusing on the robustness loss, we observe initial fluctuations during the early iterations, followed by a gradual and steady decline as the optimization progresses. By approximately the 50th iteration, the robustness loss stabilizes, indicating that the generated CFs become increasingly consistent under small perturbations of input features. This behavior demonstrates that the optimization framework effectively enhances the resilience of CF explanations, improving their reliability in real-world applications.